Forces of 70 N at 50 degrees, and 20 N at an angle
Answer #1

A force \small \overrightarrow{F_{1}} =70 N at an angle of 50o is acting on the object counterclock wise from +x - axis. So, x and y components of \small \overrightarrow{F_{1}} will be;

  \small \overrightarrow{F_{1_{_{_{_{x}}}}}}=70cos(50^{o})=45\: N

  \small \overrightarrow{F_{1_{_{_{y}}}}}=70sin(50^{o})=53.6\: N

The force \small \overrightarrow{F_{2}} =20 N at an angle of 220o is acting on the object counterclock wise from +x - axis. So, x and y components of \small \overrightarrow{F_{2}} will be;

  \small \overrightarrow{F_{2_{_{_{x}}}}}=20cos(220^{o})=-15.3\: N

  \small \overrightarrow{F_{2_{_{_{y}}}}}=20sin(220^{o})=-12.9\: N

Now by adding the x and y components of \small \overrightarrow{F_{1}} and \small \overrightarrow{F_{2}} , x and y components of resultant force \small \overrightarrow{R} can be calculated;

\small \overrightarrow{R_{x}}=\overrightarrow{F_{1_{_{_{x}}}}}+\overrightarrow{F_{2_{_{_{x}}}}}=45+(-15.3)=29.7\: N

\small \overrightarrow{R_{y}}=\overrightarrow{F_{1_{_{_{y}}}}}+\overrightarrow{F_{2_{_{_{y}}}}}=53.6+(-12.9)=40.7\: N

Therefore the magnitude of resultant force \small \overrightarrow{R} can be calculated as;

  \small \overrightarrow{R}=\sqrt{\left ( \overrightarrow{R_{x}} \right )^{2}+\left ( \overrightarrow{R_{y}} \right )^{2}}=\sqrt{(29.7)^{2}+(40.7)^{2}}

  \small \Rightarrow \overrightarrow{R}=50.38\: N

The angle of resultant force with respect to x - axis will be;

\small \theta =tan^{-1}\left (\frac{\overrightarrow{R_{y}}}{\overrightarrow{R_{x}}} \right )=tan^{-1}\left (\frac{40.7}{29.7} \right )=53.88^{o}